Determine the intersection of lines 2x+3y-15=0
and x= 2+3t
y=-6-2t

Answer by Alan3354(69443)

Determine the intersection of lines 2x+3y-15=0
and x= 2+3t
y=-6-2t
=================
x= 2+3t --> t = (x-2)/3
y=-6-2t
-------------
y=-6-2t
Sub for t
y = -6 - 2(x-2)/3
3y = -18 - 2x + 4 = -2x - 14
2x + 3y = -14
2x + 3y = 15
==========================
They're parallel, no intersection.